Solve the inequality
[4x-3]<5
In this problem the < has a line below it.

Answer by RAY100(1637)

[ 4x-3] <= 5
.
(4x-3) <= +5,,,,,,,,,,,,(4x-3) >= -5
.
4x <= 8,,,,,,,,,,,,,,,,,,4x >= -2
.
x<= 2,,,,,,,,,,,,,,,,,,,,x>= -1/2
.
check,,(x<=2),,,let x=1,,,[4(1) -3] = [1]=1 <=5,,,,ok
.
(x>= -1/2),,,let x=0,,,,,[4(0) -3] = [-3] =3 <=5,,,,ok
.